evaluate each series to the given term
-5 -25 -45 -....; 9th term

Answer by richwmiller(17219)

add -20 to last term
an = a1 + (n - 1)d
a9 = -5 + (8)(-20)
a9=-165
-5 -25 -45 -65 -85 -105 -125 -145 -165